OpenStudy (anonymous):

find the indicated power using De Moivre's Theorem (1 + i)^20

sammixboo (sammixboo):

(1 + i )^20 = { ( 1 + i )^2 } ^10 { 1 + i² + 2i } ^10 { 1 - 1 + 2i } ^10 { 2i } ^ 10 = = ( 2i x 2i x 2i x 2i x 2i ) ^2 = ( 32i^5 ) ^2 = 32i^5 x 32i^5 = 1024i^10 = 1024 x -1 = -1024

OpenStudy (anonymous):

thank you

sammixboo (sammixboo):

No problem :) Any other questions?

OpenStudy (anonymous):

This way, you did not use De Moivre Theorem

OpenStudy (anonymous):

I just have one more question. I need help to find the indicated roots, and graph the roots in the complex plane. the fifth roots of 32

sammixboo (sammixboo):

The answer to that is 2

OpenStudy (anonymous):

And what about the complex roots?

OpenStudy (anonymous):

\[ (1+i)^{20}=\left(\sqrt{2} \left(\cos \left(\frac{\pi }{4}\right)+i \sin \left(\frac{\pi }{4}\right)\right)\right)^{20}=2^{10} (\cos (5 \pi )+i \sin (5 \pi ))=-2^{10}=-1024 \]

zepdrix (zepdrix):

Mmm yah I was thinking the same thing elia +_+\[\Large\bf\sf \left(1+\mathcal i\right)^{20}\quad=\quad \left(\sqrt{2}\left[\frac{\sqrt2}{2}+\frac{\sqrt2}{2}\mathcal i\right]\right)^{20}\] \[\Large\bf\sf 2^{10}\left(\cos\frac{\pi}{4}+\mathcal i \sin\frac{\pi}{4}\right)^{20}\quad=\quad 2^{10}\left(\cos\frac{20\pi}{4}+\mathcal i\sin\frac{20\pi}{4}\right)\]

zepdrix (zepdrix):

@SithsAndGiggles Hmm I'm confused... Did you combine both questions together or something? +_+ It looks like you took the 5th root of 1+i.

OpenStudy (anonymous):

haha yeah i did... Sorry about that!

zepdrix (zepdrix):

XD

OpenStudy (anonymous):

I'll just show myself out.

OpenStudy (anonymous):

For a *general* complex number \(a+ib=r(\cos\theta+i\sin\theta)\), the \(n\)th roots are \[r^{1/n}\left(\cos\left(\frac{\theta}{n}+\frac{2k\pi}{n}\right)+i\sin\cos\left(\frac{\theta}{n}+\frac{2k\pi}{n}\right)\right)\] where \(k=0,1,2,...,n-1\).

zepdrix (zepdrix):

i sin cos( ... ) ? Oh man you're on a roll today sith XD hehe jk